### v1

# Simultaneous quantum identity authentication scheme utilizing entanglement swapping with secret key preservation

Abstract
Unconditional security in quantum key distribution (QKD) relies on authenticating the identities of users involved in key distribution. While classical identity authentication schemes were initially utilized in QKD implementations, concerns regarding their vulnerability have prompted the exploration of quantum identity authentication (QIA) protocols. In this study, we introduce a new protocol for QIA, derived from the concept of controlled secure direct quantum communication. Our proposed scheme facilitates simultaneous authentication between two users, Alice and Bob, leveraging Bell states with the assistance of a third party, Charlie. Through rigorous security analysis, we demonstrate that the proposed protocol withstands various known attacks, including impersonation, intercept and resend and impersonated fraudulent attacks. Additionally, we establish the relevance of the proposed protocol by comparing it with the existing protocols of similar type.
I Introduction
With the widespread adoption of online banking, e-commerce platforms and Internet of Things (IoT) devices, end-users have become accustomed to identity authentication procedures. The increased frequency of using such applications has underscored the importance of identity authentication schemes, which are systematic methodologies for verifying the identity of legitimate users or components/devices. These schemes are integral to virtually all cryptographic tasks, including securing communication and computation processes.
Quantum cryptography has notably revolutionized the concept of security by offering unconditional security, a feature unattainable in the classical realm. Bennett and Brassard’s seminal work in 1984 introduced an unconditionally secure protocol for quantum key distribution (QKD) [1]. Unlike classical protocols, which rely on computational complexity for security, the security of the BB84 protocol stems from the principles of quantum mechanics. Subsequent QKD protocols ([2, 3, 4, 5, 6, 7] and others [8, 9, 10, 11, 12] enable secure key distribution between a sender (Alice) and a receiver (Bob). However, before initiating any QKD protocol, Alice and Bob must authenticate each other to prevent impersonation by an eavesdropper (Eve). Therefore, identity authentication is pivotal for implementing QKD and other quantum cryptographic tasks. Initially, classical authentication schemes like the Wegman–Carter scheme [13] were mentioned as the scheme to be used in implementing the authentication step in a QKD protocol. Even in Bennett and Brassard’s pioneering work [1], it was mentioned that identity authentication will be performed by using Wegman–Carter scheme. However, classical authentication schemes lack unconditional security. Presently, commercially available QKD products and most laboratory-level implementations of quantum cryptographic tasks rely on classical authentication schemes, including post-quantum schemes. Consequently, protocols for QKD and other quantum cryptographic tasks are not truly unconditionally secure or fully quantum unless unconditionally secure schemes for quantum identity authentication (QIA) are developed. Motivated by this necessity, researchers have proposed numerous QIA schemes [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], leveraging quantum resources to achieve desired security levels. These protocols typically initiate with a pre-shared small key used for authentication. Some efforts have explored the possibility of designing schemes for device-independent QIA [26] and QIA that use homomorphic encryption with qubit rotation [27].
The first QIA scheme was introduced by Crépeau et al. [14] in 1995, utilizing oblivious transfer (OT) as a cryptographic primitive. However, subsequent work by Lo and Chau [28] demonstrated that quantum OT cannot achieve unconditional security in a two-party scenario. Consequently, Crépeau et al.’s QIA scheme lacked unconditional security. It is noteworthy that OT is not the sole cryptographic task applicable to QIA scheme design. Following Crépeau et al.’s work, a series of QIA protocols emerged [15, 16, 17, 18, 19, 20, 21, 29, 30, 22]. The schemes for QIA are built upon protocols designed for various cryptographic tasks, serving as the foundation for their development. For a comprehensive understanding of QIA and the adaptation of diverse cryptographic tasks to formulate QIA schemes, refer to the review provided in Ref. [22]. Some of these tasks were adapted from protocols for secure direct quantum communication, which involve transmitting messages using quantum resources without key generation [31, 32, 33]. Before delving into a summary of existing QIA schemes derived from modified secure direct communication schemes, it is pertinent to distinguish between two types of secure quantum direct communication: quantum secure direct communication (QSDC) and deterministic secure quantum communication (DSQC). Numerous QSDC and DSQC schemes exist [31, 34, 35, 36, 37, 38, 39, 40], with many early ones being entangled state-based [31, 38, 41]. For instance, Zhang et al. [17] adapted the entangled state-based ping-pong protocol [41] for QSDC to devise a QIA scheme. Moreover, Yuan et al. [42] proposed a QIA protocol based on the LM05 protocol [32], which is the single photon analog of the ping-pong protocol. Other authors have also modified DSQC and QSDC protocols to create QIA schemes [15, 43, 44]. Additionally, novel schemes have emerged for controlled DSQC (CDSQC) [35, 45], where a semi-honest Alice securely communicates with Bob using quantum resources only under the supervision of a controller, Charlie. The potential of adapting CDSQC schemes for QIA remains relatively unexplored. Taking into account the significance of identity authentication, we aim to propose a new QIA protocol with simultaneous authentication of legitimate parties, inspired by CDSQC principles. In the following sections, we will introduce a Bell state-based protocol for QIA, drawing from CDSQC concepts, and demonstrate its resilience against common attacks.
The remainder of this letter is structured as follows: Section II introduces the proposed QIA protocol that uses Bell states. Section III critically analyzes the protocol’s security, demonstrating its robustness against multiple types of attacks. Section IV compares our protocol with the existing ones from the same family. Finally, Section V concludes the letter.
II New quantum identity authentication protocol
In this section, we aim to introduce a new framework for QIA, akin to schemes inspired by QSDC. As it utilizes Bell states, our protocol can be construed as an entangled state-driven approach to QIA. Within this framework, two legitimate parties mutually authenticate one another with the assistance of an untrusted third party, employing unitary operations. Prior to delving into the specifics of our protocol, it is important to provide a succinct overview of the underlying conceptual framework driving its design.
II.1 Principal concept
This protocol relies on the principles of Bell states with entanglement swapping and the application of Pauli operations. The corresponding connections between the Bell state and the pre-shared authentication key can be delineated as follows:
$\begin{array}[]{lcl}00:|\phi^{+}\rangle&=&\frac{1}{\sqrt{2}}\left(\left|00\right\rangle+\left|11\right\rangle\right),\\ 01:|\phi^{-}\rangle&=&\frac{1}{\sqrt{2}}\left(\left|00\right\rangle-\left|11\right\rangle\right),\\ 10:|\psi^{+}\rangle&=&\frac{1}{\sqrt{2}}\left(\left|01\right\rangle+\left|10\right\rangle\right),\\ 11:|\psi^{-}\rangle&=&\frac{1}{\sqrt{2}}\left(\left|01\right\rangle-\left|10\right\rangle\right).\end{array}$ (1)
The Eqs. (1) and (2) describe the relationship between the pre-shared key, and the Bell state prepared by Alice and Bob, as well as the Pauli operations they perform. The association between the respective Pauli operations and the pre-shared authentication key can be stated as follows:
$\begin{array}[]{lcl}00:\mathds{1}_{2}&=&\left|0\right\rangle\left\langle 0\right|+\left|1\right\rangle\left\langle 1\right|,\\ 01:\sigma_{x}&=&\left|0\right\rangle\left\langle 1\right|+\left|1\right\rangle\left\langle 0\right|,\\ 10:i\sigma_{y}&=&\left|0\right\rangle\left\langle 1\right|-\left|1\right\rangle\left\langle 0\right|,\\ 11:\sigma_{z}&=&\left|0\right\rangle\left\langle 0\right|-\left|1\right\rangle\left\langle 1\right|.\end{array}$ (2)
Let us provide a concise overview of the protocol described. Alice and Bob possess a pre-shared authentication key sequence consisting of $n+1$ pairs of bits. Each key pair represents two bits of information. Initially, let us consider a specific example where we denote the first two secret keys of the sequence as $11$ and $00$. Consider that Alice selects the second key ($00$), generally referred to as the $\left(m+1\right)^{{\rm th}}$ key in the entire sequence, while Bob selects the XOR of the second key and the first key $\left(00\oplus 11=11\right)$, which can be understood as the XOR of the $\left(m+1\right)^{{\rm th}}$ key and the $m^{{\rm th}}$ key, where $m=1,2,\cdots,n.$ Subsequently, they generate Bell states based on their selected keys using the Eq. (1). In this scenario, Alice and Bob prepare the states $\left|\phi^{+}\right\rangle_{12}$ and $\left|\psi^{-}\right\rangle_{34}$, where the subscripts $1,2$ and $3,4$ represent the particles of Alice and Bob, respectively. Similarly, they generate a series of Bell states using the pre-shared authentication key sequence. Alice retains the sequence of particle $1$ of the Bell states and transmits the sequence associated with particle $2$ of the Bell states to an untrusted third party, Charlie. Charlie applies a permutation operator $\Pi_{n}$ [36, 46, 47, 45] to the sequence of particle $2$ while keeping the original sequence confidential. Following the permutation, Charlie forwards the sequence to Bob. Bob independently prepares two sequences associated with particle $3$ and particle $4$. Bob sends the sequence corresponding to particle $4$ directly to Alice. Consequently, Alice and Bob possess the sequences containing particles $1,4$ and $2,3$, respectively. With the specific two secret keys in place, the composite system is characterized by,
$\begin{array}[]{lcl}\left|\phi^{+}\right\rangle_{12}\left|\psi^{-}\right\rangle_{34}&=&\frac{1}{2}\left(\left|\psi^{+}\right\rangle_{14}\left|\phi^{-}\right\rangle_{23}+\left|\psi^{-}\right\rangle_{14}\left|\phi^{+}\right\rangle_{23}-\left|\phi^{+}\right\rangle_{14}\left|\psi^{-}\right\rangle_{23}-\left|\phi^{-}\right\rangle_{14}\left|\psi^{+}\right\rangle_{23}\right).\end{array}$
Alice (Bob) applies a Pauli operation $\sigma_{z}$ on qubit $1$ (qubit $3$) in accordance with the first key, here considered as $11$ This [^1] For a long pre-shared key sequence, the Pauli operation will be determined by the $m^{{\rm th}}$ key. is done utilizing the map described in Eq. (2). Subsequently, Alice and Bob publicly announce the completion of their operations on the designated particles. Charlie then publicly announces the permutation operation ($\Pi_{n}$). Bob performs the inverse permutation operation on his sequence of particle 2, thereby restoring the original sequence. Following this, they conduct a Bell measurement on the particles currently in their possession (Alice performs Bell measurement on particles $1$ and $4$, while Bob on particles $2$ and $3$). The resultant composite system is as follows,
$\begin{array}[]{lcl}\left|\Psi^{11}\right\rangle&=&\sigma_{z1}\otimes\sigma_{z3}\left(\left|\phi^{+}\right\rangle_{12}\left|\psi^{-}\right\rangle_{34}\right)\\ &=&\frac{1}{2}\left(\left|\psi^{-}\right\rangle_{14}\left|\phi^{+}\right\rangle_{23}+\left|\psi^{+}\right\rangle_{14}\left|\phi^{-}\right\rangle_{23}-\left|\phi^{-}\right\rangle_{14}\left|\psi^{+}\right\rangle_{23}-\left|\phi^{+}\right\rangle_{14}\left|\psi^{-}\right\rangle_{23}\right).\end{array}$ (3)
Alice (Bob) converts the outcome of her (his) Bell measurement into classical bits using Eq. (1). The measurement result is not disclosed. In our scenario, Alice has measurement outcomes, each with a probability of $\frac{1}{4}$, represented as $\left|\psi^{-}\right\rangle,$ $\left|\psi^{+}\right\rangle,$ $\left|\phi^{-}\right\rangle$ and $\left|\phi^{+}\right\rangle$, with corresponding classical bit values of 11, 10, 01 and 00. The measurement outcomes of Bob correspond to $\left|\phi^{+}\right\rangle,$ $\left|\phi^{-}\right\rangle,$ $\left|\psi^{+}\right\rangle$ and $\left|\psi^{-}\right\rangle$, with corresponding classical bit values 00, 01, 10 and 11, respectively. In this context, Alice assumes the role of the verifier since we are considering two secret keys in our example. Suppose Bob declares his bit value as 00 through an unjammable classical communication channel. Then Alice performs an XOR operation on Bob’s announcement and her classical bit value (11) that corresponds to her measurement outcome. If no errors occur in the quantum channel due to noise or eavesdropping, the XOR value should match the first key of the initial two secret key sequences, which is 11. It is important to note that generally, the XOR value would correspond to the $m^{{\rm th}}$ key when considering the pre-shared key sequence. The remaining possible outcomes and their correlations are detailed in Table 1 when the initial pre-shared keys are 11 and 00.
Alice and Bob’s possible measurement outcomes Additional modulo 2
$\left|\psi^{-}\right\rangle_{14}\otimes\left|\phi^{+}\right\rangle_{23}$ $11\oplus 00=11$
$\left|\psi^{+}\right\rangle_{14}\otimes\left|\phi^{-}\right\rangle_{23}$ $10\oplus 01=11$
$\left|\phi^{-}\right\rangle_{14}\otimes\left|\psi^{+}\right\rangle_{23}$ $01\oplus 10=11$
$\left|\phi^{+}\right\rangle_{14}\otimes\left|\psi^{-}\right\rangle_{23}$ $00\oplus 11=11$
Table 1: The possible measurement results of the legitimate parties for our QIA scheme.
We take 3 more examples to ensure the persistence of our conclusion across the remaining three possible secret keys in the $m^{{\rm th}}$ position of the pre-shared key sequence, namely {00}, {01} and {10}, we can depict the final state shared by Alice and Bob. The final composite systems, considering the secret keys at positions $m^{{\rm th}}$ and $\left(m+1\right)^{{\rm th}}$ of the entire sequence to be $\left\{00,01\right\},$$\left\{01,10\right\}$ and $\left\{10,11\right\}$ respectively, are expressed as follows:
$\begin{array}[]{lcl}\left|\Psi^{00}\right\rangle&=&\mathds{1}_{1}\otimes\mathds{1}_{3}\left(\left|\phi^{-}\right\rangle_{12}\left|\phi^{-}\right\rangle_{34}\right)\\ &=&\frac{1}{2}\left(\left|\phi^{+}\right\rangle_{14}\left|\phi^{+}\right\rangle_{23}+\left|\phi^{-}\right\rangle_{14}\left|\phi^{-}\right\rangle_{23}-\left|\psi^{+}\right\rangle_{14}\left|\psi^{+}\right\rangle_{23}-\left|\psi^{-}\right\rangle_{14}\left|\psi^{-}\right\rangle_{23}\right)\end{array},$ (4)
$\begin{array}[]{lcl}\left|\Psi^{01}\right\rangle&=&\sigma_{x1}\otimes\sigma_{x3}\left(\left|\psi^{+}\right\rangle_{12}\left|\psi^{-}\right\rangle_{34}\right)\\ &=&\frac{1}{2}\left(-\left|\phi^{+}\right\rangle_{14}\left|\phi^{-}\right\rangle_{23}-\left|\phi^{-}\right\rangle_{14}\left|\phi^{+}\right\rangle_{23}+\left|\psi^{+}\right\rangle_{14}\left|\psi^{-}\right\rangle_{23}+\left|\psi^{-}\right\rangle_{14}\left|\psi^{+}\right\rangle_{23}\right)\end{array},$ (5)
and
$\begin{array}[]{lcl}\left|\Psi^{10}\right\rangle&=&i\sigma_{y1}\otimes i\sigma_{y3}\left(\left|\psi^{-}\right\rangle_{12}\left|\phi^{-}\right\rangle_{34}\right)\\ &=&\frac{1}{2}\left(\left|\psi^{+}\right\rangle_{14}\left|\phi^{+}\right\rangle_{23}+\left|\psi^{-}\right\rangle_{14}\left|\phi^{-}\right\rangle_{23}+\left|\phi^{+}\right\rangle_{14}\left|\psi^{+}\right\rangle_{23}+\left|\phi^{-}\right\rangle_{14}\left|\psi^{-}\right\rangle_{23}\right)\end{array},$ (6)
respectively. In Eqs. (3 - 6), we consider 4 combinations of $m^{{\rm th}}$ and $\left(m+1\right)^{{\rm th}}$ terms. However, there are 12 additional combinations that yield the same conclusive measurement results which provide authentication for legitimate parties.
II.2 Protocol description
In the scenario considered in this letter, two legitimate parties, Alice and Bob, seek to authenticate themselves with the help of an untrusted party Charlie. Within this authentication protocol, Alice and Bob possess a pre-shared classical secret key denoted as $K_{AB}=\left\{k_{1}^{1}k_{2}^{1},k_{1}^{2}k_{2}^{2},k_{1}^{3}k_{2}^{3},\cdots,k_{1}^{m}k_{2}^{m},\cdots k_{1}^{n+1}k_{2}^{n+1}\right\}$. The secret key sequence $K_{AB}$ comprises two bit key information with uniform distribution, denoted by $k_{1}^{m}k_{2}^{m}$, where $k_{1}^{m}k_{2}^{m}\in\{00,01,10,11\}$. Before delving into the protocol’s procedural details, it is pertinent to briefly elucidate the utility of decoy states in establishing secure communication channels between any two parties [48]. Typically, one party randomly inserts decoy qubits, ideally equal in number to the original information sequence, into the information qubit sequence. This enlarged sequence is then transmitted to the other party. Upon receiving this enlarged sequence, the receiving party asks for positional and encoding details of the decoy qubits from the transmitting party via an impervious public channel. The transmitting party declares this requisite information publicly, enabling the receiving party to verify it by measuring the decoy qubits. Should errors exceeding the acceptable threshold be detected, both parties terminate the entire qubit sequence. This utilization of decoy states heightens the security of the channel and facilitates the secure transmission of information between parties. Now, let’s delineate the sequential steps involved in the authentication process:
Step 1
Alice and Bob generate Bell state sequences $A_{12}$ and $B_{34}$ using the $\left(m+1\right)^{{\rm th}}$ key and XOR operation of the $m^{{\rm th}}$ and $\left(m+1\right)^{{\rm th}}$ keys of the entire sequence $K_{AB}$, respectively, as stipulated by the procedure outlined in Eq. (1), where $m=1,2,\cdots,n$.
$\begin{array}[]{lcl}A&=&\left\{|A\rangle_{12}^{1},|A\rangle_{12}^{2},|A\rangle_{12}^{3},\cdots,|A\rangle_{12}^{m},\cdots,|A\rangle_{12}^{n}\right\},\\ B&=&\left\{|B\rangle_{34}^{1},|B\rangle_{34}^{2},|B\rangle_{34}^{3},\cdots,|B\rangle_{34}^{m},\cdots,|B\rangle_{34}^{n}\right\}.\end{array}$
The subscripts $1,2$ and $3,4$ correspond to the particles assigned to Alice and Bob, respectively. Ideally, these two sets should be identical.
Step 2
Alice and Bob partition the states of their Bell pairs into two distinct sequences of $n$ particles. In each sequence, the first particle of every Bell pair constitutes one sequence, while the second particle of each Bell pair forms the other sequence. Thus, Alice and Bob possess the sequences denoted as $S_{A1},\,S_{A2}$ and $S_{B3},\,S_{B4}$, respectively.
$\begin{array}[]{lcl}S_{A1}&=&\left\{s_{1}^{1},s_{1}^{2},s_{1}^{3},\cdots,s_{1}^{m},\cdots,s_{1}^{n}\right\},\\ S_{A2}&=&\left\{s_{2}^{1},s_{2}^{2},s_{2}^{3},\cdots,s_{2}^{m},\cdots,s_{2}^{n}\right\},\\ S_{B3}&=&\left\{s_{3}^{1},s_{3}^{2},s_{3}^{3},\cdots,s_{3}^{m},\cdots,s_{3}^{n}\right\},\\ S_{B4}&=&\left\{s_{4}^{1},s_{4}^{2},s_{4}^{3},\cdots,s_{4}^{m},\cdots,s_{4}^{n}\right\}.\end{array}$
Here, $S_{A1}$ ($S_{A2}$) represents the sequence comprising the first (second) particles of all Bell states within sequence $A$. Correspondingly, $S_{B3}$ ($S_{B4}$) denotes the sequence encompassing the first (second) particles of all Bell states within sequence $B$. Alice and Bob possess the particle sequences $S_{A1}$ and $S_{B3}$ with themselves, respectively. Additionally, Alice (Bob) procedurally introduces the decoy particles $D_{A}$ $\left(D_{B}\right)$ into the sequence $S_{A2}$ ($S_{B4}$) to formulate an augmented sequence $S_{A2}^{\prime}$ ($S_{B4}^{\prime}$). Alice (Bob) subsequently transmits the sequence $S_{A2}^{\prime}$ ($S_{B4}^{\prime}$) via the quantum channel to Charlie (Alice).
Step 3
Charlie receives the sequence $S_{A2}^{\prime}$ and performs security tests utilizing decoy particles. Following the successful completion of the security tests, Charlie eliminates the decoy particles $D_{A}$ to retrieve the original sequence $S_{A2}$. Upon obtaining $S_{A2}$, Charlie applies a permutation operation $\Pi_{n}$ to the sequence and introduces decoy states to produce the new sequence $S_{A2}^{*}$. Subsequently, Charlie transmits the new sequence $S_{A2}^{*}$ to Bob.
Step 4
Alice and Bob perform a security test employing decoy particles. Upon the successful conclusion of the security assessment, Alice eliminates the decoy states denoted as $D_{B}$ from the enlarged sequence $S_{B4}^{{}^{\prime}}$, reverting it to the initial sequence prepared by Bob, designated as $S_{B4}$. Subsequently, Charlie publicly discloses the precise ordering of the sequence $S_{A2}^{*}$ which is a reverse operation of $\Pi_{n}$ operator. Bob reorganizes the sequence to restore its original configuration, denoted as $S_{A2}$.
Step 5
Alice (Bob) applies Pauli operations to all $m^{{\rm th}}$ particles in sequence $S_{A1}$ ($S_{B3}$) based on the $m^{{\rm th}}$ key of the pre-shared key sequence, utilizing Eq. (2). Subsequently, Alice (Bob) possesses sequences $S_{A1},\,S_{B4}$ ($S_{A2},\,S_{B3}$) and conducts Bell measurements on the respective particles of these sequences [^2] Perform Bell measurement on $m^{{\rm th}}$ particle of sequence $S_{A1}$($S_{A2}$) and $m^{{\rm th}}$ particle of sequence $S_{B4}$ ($S_{B3}$).. Alice and Bob record the measurement outcomes of the Bell state measurement as classical key sequences $R_{14}=\left\{r_{14}^{1},r_{14}^{2},r_{14}^{3},\cdots,r_{14}^{m},\cdots,r_{14}^{n}\right\}$ and $R_{23}=\left\{r_{23}^{1},r_{23}^{2},r_{23}^{3},\cdots,r_{23}^{m},\cdots,r_{23}^{n}\right\}$ following Eq. (1).
Step 6
Alice publicly discloses the position and values of $\frac{n}{2}$ keys from the sequence $R_{14}$, while Bob also discloses $\frac{n}{2}$ keys from the sequence $R_{23}$ of corresponding remaining position of the sequence $R_{14}$. Alice (Bob) executes XOR operations between Bob’s (Alice’s) announced key and the corresponding keys from her (his) sequence $R_{14}$ ($R_{23}$), resulting in a new two-bit key sequence $r_{A}$ ($r_{B}$) of length $n$.
Step 7
In an error-free scenario, the elements of $r_{A}$ and $r_{B}$ should match with the corresponding key elements of the initial sequence $K_{AB}$ [^3] Here, $n$ elements of sequences of $r_{A}$ and $r_{B}$ should match with the corresponding $n$ elements of sequence of $K_{AB}$. It may be noted that $K_{AB}$ has $n+1$ pre-shared keys.. Upon meeting this criterion, authentication is successfully attained for both Alice and Bob. This authentication process is executed simultaneously by the two legitimate parties.
The presented QIA protocol is depicted in a flowchart illustrated in Fig. 1.
Figure 1: Flowchart illustrating the operation of the proposed QIA protocol.
III Evaluation of the security aspects in the proposed protocol
In this section, we aim to evaluate the security of the proposed protocol against various well-known attacks that can be attempted by an eavesdropper, commonly referred to as Eve. It is important to mention that insider attacks (by Alice and Bob) will not be discussed here, as they are not pertinent within the scope of QIA. Initially, we will focus on the impersonation attack. In this scenario, we consider the possibility of an untrusted participant, Charlie, attempting to gain unauthorized access to confidential information while adhering to all steps of the protocol. Moreover, an impersonated fraudulent attack entails an external party, such as Eve, aiming to mimic the legitimate user, either Alice or Bob, and effectively maneuver through the authentication procedure.
III.1 Security assessment of the protocol against impersonation attack by Eve
Without loss of generality, let us consider the scenario outlined in Section II.1. Eve, acting as Alice, selects the state $|\phi^{+}\rangle_{12}$ and transmits the particle (particle 2) sequence $S_{e2}$ to Charlie [^4] We adopt the term "sequence" for the sake of generality. However, to elucidate the intricacies of this security analysis, we employ an illustrative example outlined in Section II.1.. Charlie adheres to the protocol by applying the permutation operation and forwards the sequence $S_{e2}^{\prime}$ to Bob. Upon receiving Charlie’s announcement regarding the correct sequence order of $S_{e2}^{\prime}$, Bob retrieves the original $S_{e2}$. Additionally, Bob sends the particle sequence $S_{B4}$ directly to Eve. Eve, while masquerading as Alice, executes all protocol steps as expected, applying the Pauli operation (here $\sigma_{z1}$) to Particle $1$. Similarly, Bob performs the Pauli operation (here $\sigma_{z3}$) on his particle 3. Subsequently, Bob and Eve conduct Bell measurements, and Table (1) illustrates all possible measurement scenarios revealing Eve’s presence.
As Eve does not possess knowledge of the correct key pair $K_{AB}$ (e.g., 11 and 00) nor the permutation operation performed by Charlie, $\Pi_{n}$, her probability of selecting the correct Bell state is $\frac{1}{4}$. To successfully evade detection, Eve must accurately guess all $n$ Bell states correctly. Hence, the probability of Eve executing a successful impersonation attack is $\left(\frac{1}{4}\right)^{n}$. As $n$ grows large, the likelihood of a successful impersonation attack diminishes nearly to zero. Consequently, the probability $P\left(n\right)$ of detecting Eve’s presence is $1-\left(\frac{1}{4}\right)^{n}$. For sufficiently large $n$ values, $P\left(n\right)$ approximates $1$, facilitating the identification of an impersonation attack without failure. The relationship between $P\left(n\right)$ and $n$ is depicted in Fig. 2, which illustrates the necessity of at least 6 pre-shared keys to detect Eve’s presence effectively.
Figure 2: The correlation between the probability $P\left(n\right)$ of detecting Eve’s presence and the number of classical keys $n$ used as a pre-shared authentication key.
III.2 Security assessment of the protocol against intercept and resend attack
In this quantum protocol, Alice and Bob generate Bell states based on their pre-shared secret key. Instead of transmitting transmitting their entire quantum states, Alice and Bob selectively send Particle 2 and Particle 4 through the quantum channel. This approach limits Eve’s potential attacks to these specific particles at any given moment. For the sake of simplicity, we assume that Eve will attempt to intercept Particle 2 and Particle 4 as they are transmitted from Alice to Charlie and from Bob to Alice, respectively. Eve’s ability to extract information from the quantum channel is constrained by the Holevo bound or Holevo quantity [49]
[…]
where $\rho^{24}$ and $\rho_{i}^{24}$ represent the reduced matrices of $\rho$ and $\rho_{i}$, respectively, obtained after performing a partial trace over particles 1 and 3. Through straightforward calculations, we determine that $\rho^{24}={\rm Tr}_{13}\left(\rho\right)=\frac{1}{4}\left(|11\rangle\langle 11|+|10\rangle\langle 10|+|01\rangle\langle 01|+|00\rangle\langle 00|\right)=\frac{1}{4}\mathds{1}_{4}$, where $\mathds{1}_{4}$ denotes the $4\times 4$ identity matrix. Similarly, the von Neumann entropy of each component of the mixed state $\rho$ after the partial trace, denoted as $\rho_{i}^{24}={\rm Tr}_{13}\left(\rho_{i}\right)=\frac{1}{4}\mathds{1}_{4}$. By substituting $\rho^{24}$ and $\rho_{i}^{24}$ into Eq. (8), we deduce that $\chi\left(\rho^{24}\right)=0$. Consequently, we infer that Eve cannot obtain any key information through direct intercept attacks on the transmission particles. The intercept-resend (IR) attack strategy employed by Eve is depicted in Fig. 3.
Figure 3: Eve’s IR attack strategy. Here, the numbers 1, 2, 3 and 4 represent the qubits in sequences $S_{A1}$, $S_{A2}$, $S_{B3}$ and $S_{B4}$, respectively.
III.3 Security assessment of the protocol against impersonated fraudulent attack
To demonstrate our protocol against impersonated fraudulent attack, we consider two pre-shared secret keys as 11 and 00, respectively. Initially, Alice and Bob prepare the Bell state $|\phi^{+}\rangle_{12}$ and $|\psi^{-}\rangle_{34}$, respectively. Meanwhile, Eve prepares two new single-qubit states $|\chi\rangle_{5}$ and $|\chi\rangle_{6}$ as fake states, which are described as $a|0\rangle+b|1\rangle$ and $c|0\rangle+d|1\rangle$, respectively, where $\left|a\right|^{2}+\left|b\right|^{2}=1$ and $\left|c\right|^{2}+\left|d\right|^{2}=1$. Eve then performs a CNOT operation with particle 2 (4) acting as the control qubit and the target qubit being particle 5 (6). She retains particles 2 and 4, and sends particles 5 and 6 to Bob and Alice, respectively. The composite state following Eve’s attack, as induced by the CNOT operation, is expressed as follows:
$\begin{array}[]{lcl}&&{\rm CNOT_{2(4)\rightarrow 5(6)}|\phi^{+}\rangle_{12}\otimes|\psi^{-}\rangle_{34}\otimes|\chi\rangle_{5}\otimes|\chi\rangle_{6}}\\ &=&\frac{1}{2}\left[|0001\rangle\left(a|0\rangle+b|1\rangle\right)\left(c|1\rangle+d|0\rangle\right)-|0010\rangle\left(a|0\rangle+b|1\rangle\right)\left(c|0\rangle+d|1\rangle\right)\right.\\ &+&\left.|1101\rangle\left(a|1\rangle+b|0\rangle\right)\left(c|1\rangle+d|0\rangle\right)-|1110\rangle\left(a|1\rangle+b|0\rangle\right)\left(c|0\rangle+d|1\rangle\right)\right]_{123456}\\ &=&\frac{1}{2}\left[ac|000101\rangle+ad|000100\rangle+bc|000111\rangle+bd|000110\rangle\right.\\ &-&ac|001000\rangle-ad|001001\rangle-bc|001010\rangle-bd|001011\rangle\\ &+&ac|110111\rangle-ad|110110\rangle-bc|110101\rangle-bd|110100\rangle\\ &-&\left.ac|111010\rangle-ad|111011\rangle-bc|111000\rangle-bd|111001\rangle\right]_{123456}\end{array}.$ (9)
As per the protocol, Alice and Bob execute $\sigma_{z}$ operation on their respective particles 1 and 3. Subsequently, they conduct Bell measurements on particles 1 and 6 for Alice, and on particles 5 and 3 for Bob. Eve is assigned particles 2 and 4. The resultant composite system of Alice, Bob and Eve, subsequent to the Pauli operation and Bell measurement described in Eq. (9), can be represented as:
$\begin{array}[]{lcl}|\Psi\rangle&=&\frac{1}{2\sqrt{2}}\left[ac\left(|\psi^{+}\rangle|\phi^{+}\rangle|\psi^{-}\rangle+|\psi^{+}\rangle|\phi^{-}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\phi^{-}\rangle|\psi^{+}\rangle\right.\right.\\ &+&\left.|\phi^{+}\rangle|\psi^{+}\rangle|\phi^{-}\rangle+|\phi^{+}\rangle|\psi^{-}\rangle|\phi^{+}\rangle+|\phi^{-}\rangle|\psi^{+}\rangle|\phi^{+}\rangle+|\phi^{-}\rangle|\psi^{-}\rangle|\phi^{-}\rangle\right)\\ &+&ad\left(|\phi^{+}\rangle|\phi^{+}\rangle|\psi^{-}\rangle+|\phi^{+}\rangle|\phi^{-}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\phi^{+}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\phi^{-}\rangle|\psi^{-}\rangle\right.\\ &+&\left.|\psi^{+}\rangle|\psi^{+}\rangle|\phi^{-}\rangle+|\psi^{+}\rangle|\psi^{-}\rangle|\phi^{+}\rangle+|\psi^{-}\rangle|\psi^{+}\rangle|\phi^{+}\rangle+|\psi^{-}\rangle|\psi^{-}\rangle|\phi^{-}\rangle\right)\\ &+&bc\left(|\psi^{+}\rangle|\psi^{+}\rangle|\psi^{-}\rangle-|\psi^{+}\rangle|\psi^{-}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\psi^{+}\rangle|\psi^{+}\rangle-|\psi^{-}\rangle|\psi^{-}\rangle|\psi^{-}\rangle\right.\\ &+&\left.|\phi^{+}\rangle|\phi^{+}\rangle|\phi^{-}\rangle-|\phi^{+}\rangle|\phi^{-}\rangle|\phi^{+}\rangle+|\phi^{-}\rangle|\phi^{+}\rangle|\phi^{+}\rangle-|\phi^{-}\rangle|\phi^{-}\rangle|\phi^{-}\rangle\right)\\ &&bd\left(|\phi^{+}\rangle|\psi^{+}\rangle|\psi^{-}\rangle-|\phi^{+}\rangle|\psi^{-}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\psi^{+}\rangle|\psi^{+}\rangle-|\phi^{-}\rangle|\psi^{-}\rangle|\psi^{-}\rangle\right.\\ &+&\left.\left.|\psi^{+}\rangle|\phi^{+}\rangle|\phi^{-}\rangle-|\psi^{+}\rangle|\phi^{-}\rangle|\phi^{+}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle|\phi^{+}\rangle-|\psi^{-}\rangle|\phi^{-}\rangle|\phi^{-}\rangle\right)\right]_{165324}\end{array}.$ (10)
When Alice and Bob obtain Bell pairs in the states $\left|\psi^{+}\right\rangle_{16}\left|\phi^{-}\right\rangle_{53},\left|\psi^{-}\right\rangle_{16}\left|\phi^{+}\right\rangle_{53},\left|\phi^{+}\right\rangle_{16}\left|\psi^{-}\right\rangle_{53}\,{\rm and}\left|\phi^{-}\right\rangle_{16}\left|\psi^{+}\right\rangle_{53}$ after measurement, detection probability of eavesdropper’s presence becomes zero. By analyzing the density operator of the final composite state $|\Psi\rangle\langle\Psi|$, as from Eq. (10), we can determine the non-detection probability of Eve’s presence, denoted as ${\rm P_{nd}}$, to be $\frac{1}{2}\left(\left|ac\right|^{2}+\left|bd\right|^{2}\right)$. Therefore, the detection probability of Eve’s presence, denoted as ${\rm P_{d}}$, is ${\rm P_{d}=1-}\frac{1}{2}\left(\left|ac\right|^{2}+\left|bd\right|^{2}\right)$. Eve can minimize her detection probability by setting $\left|a\right|=\left|b\right|=\left|c\right|=\left|d\right|=\frac{1}{\sqrt{2}}$. Thus, the single qubit states of Eve are $|+\rangle_{5}=\frac{1}{\sqrt{2}}\left(|0\rangle+|1\rangle\right)_{5}$ and $|+\rangle_{6}=\frac{1}{\sqrt{2}}\left(|0\rangle+|1\rangle\right)_{6}$. Under this condition, ${\rm P_{d}}=\frac{3}{4}$, representing the minimum detection probability of Eve when she selects the single qubit state as a fake state.
Now, we investigate the same scenario with the only difference being Eve’s entangled state as her fake state to impersonate herself as a legitimate party. In this scenario, we consider the same pair of pre-shared keys, $11,00$, and Eve’s fake state is $|\chi^{\prime}\rangle_{56}=\left(a^{\prime}|00\rangle+b^{\prime}|01\rangle+c^{\prime}|10\rangle+d^{\prime}|11\rangle\right)_{56}$, where $\left|a^{\prime}\right|^{2}+\left|b^{\prime}\right|^{2}+\left|c^{\prime}\right|^{2}+\left|d^{\prime}\right|^{2}=1$. The remaining process is the same as the previous one. Eventually, the final composite system of Alice, Bob and Eve after their Bell measurement is
$\begin{array}[]{lcl}|\Psi^{\prime}\rangle&=&\frac{1}{2\sqrt{2}}\left[a^{\prime}\left(|\psi^{+}\rangle|\phi^{+}\rangle|\psi^{-}\rangle+|\psi^{+}\rangle|\phi^{-}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\phi^{-}\rangle|\psi^{+}\rangle\right.\right.\\ &+&\left.|\phi^{+}\rangle|\psi^{+}\rangle|\phi^{-}\rangle+|\phi^{+}\rangle|\psi^{-}\rangle|\phi^{+}\rangle+|\phi^{-}\rangle|\psi^{+}\rangle|\phi^{+}\rangle+|\phi^{-}\rangle|\psi^{-}\rangle|\phi^{-}\rangle\right)\\ &+&b^{\prime}\left(|\phi^{+}\rangle|\phi^{+}\rangle|\psi^{-}\rangle+|\phi^{+}\rangle|\phi^{-}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\phi^{+}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\phi^{-}\rangle|\psi^{-}\rangle\right.\\ &+&\left.|\psi^{+}\rangle|\psi^{+}\rangle|\phi^{-}\rangle+|\psi^{+}\rangle|\psi^{-}\rangle|\phi^{+}\rangle+|\psi^{-}\rangle|\psi^{+}\rangle|\phi^{+}\rangle+|\psi^{-}\rangle|\psi^{-}\rangle|\phi^{-}\rangle\right)\\ &+&c^{\prime}\left(|\psi^{+}\rangle|\psi^{+}\rangle|\psi^{-}\rangle-|\psi^{+}\rangle|\psi^{-}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\psi^{+}\rangle|\psi^{+}\rangle-|\psi^{-}\rangle|\psi^{-}\rangle|\psi^{-}\rangle\right.\\ &+&\left.|\phi^{+}\rangle|\phi^{+}\rangle|\phi^{-}\rangle-|\phi^{+}\rangle|\phi^{-}\rangle|\phi^{+}\rangle+|\phi^{-}\rangle|\phi^{+}\rangle|\phi^{+}\rangle-|\phi^{-}\rangle|\phi^{-}\rangle|\phi^{-}\rangle\right)\\ &&d^{\prime}\left(|\phi^{+}\rangle|\psi^{+}\rangle|\psi^{-}\rangle-|\phi^{+}\rangle|\psi^{-}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\psi^{+}\rangle|\psi^{+}\rangle-|\phi^{-}\rangle|\psi^{-}\rangle|\psi^{-}\rangle\right.\\ &+&\left.\left.|\psi^{+}\rangle|\phi^{+}\rangle|\phi^{-}\rangle-|\psi^{+}\rangle|\phi^{-}\rangle|\phi^{+}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle|\phi^{+}\rangle-|\psi^{-}\rangle|\phi^{-}\rangle|\phi^{-}\rangle\right)\right]_{165324}\end{array}.$ (11)
In the scenario described, the detection probability of Eve’s presence is null when Alice and Bob obtain measurement results associated with specific Bell pairs: $\left|\psi^{+}\right\rangle_{16}\left|\phi^{-}\right\rangle_{53},\left|\psi^{-}\right\rangle_{16}\left|\phi^{+}\right\rangle_{53},\left|\phi^{+}\right\rangle_{16}\left|\psi^{-}\right\rangle_{53}\,{\rm and}\left|\phi^{-}\right\rangle_{16}\left|\psi^{+}\right\rangle_{53}$. The non-detection probability of Eve’s presence can be computed from the density matrix, $|\Psi^{\prime}\rangle\langle\Psi^{\prime}|$ from Eq. (11), denoted as ${\rm P_{nd}}=\frac{1}{2}\left(\left|a^{\prime}\right|^{2}+\left|d^{\prime}\right|^{2}\right)$. Correspondingly, the detection probability of Eve’s presence, denoted as ${\rm P_{d}}=1-\frac{1}{2}\left(\left|a^{\prime}\right|^{2}+\left|d^{\prime}\right|^{2}\right)$. Eve strategically sets $\left|a^{\prime}\right|=\left|d^{\prime}\right|=\frac{1}{\sqrt{2}}$ and $\left|b^{\prime}\right|=\left|c^{\prime}\right|=0$ to minimize her detection probability. The entangled state of Eve is represented by $|\phi^{+}\rangle_{56}=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)_{56}$, resulting in ${\rm P_{d}}=\frac{1}{2}$. It follows that using an entangled state rather than a single qubit state reduces the detection probability of Eve’s presence. Consequently, the minimum detection probabilities for single qubit and entangled states are $\frac{3}{4}$ and $\frac{1}{2}$ respectively. Based on the discussion, it can be deduced that the protocol is secure against impersonated fraudulent attack by Eve, particularly under optimal conditions for Eve.
IV Comparison with the existing protocols
In this section, we aim to conduct a concise comparative analysis involving the proposed protocol and a selection of previously proposed QIA schemes. The comparison focuses on several key aspects: the utilization of quantum resources, the nature of the third-party involvement (whether they are assumed honest/semi-honest or considered untrusted), the minimum number of pre-shared keys required for authentication, and whether the scheme enables bidirectional mutual authentication or only allows unidirectional authentication where it permits only one legitimate party to check the authenticity of the other. Given the abundance of QIA schemes, we have carefully chosen representative ones for comparison, with particular attention to schemes based on protocols for secure direct quantum communication, akin to the present protocol, which draws inspiration from CDSQC.
We commence the comparison by examining Zhang et al.’s QIA scheme [17], which relied on the ping-pong protocol for QSDC. In Zhang et al.’s protocol, Alice serves as the reliable certification authority, and Bob represents the common user whose identity requires verification by Alice. Similarly, Yuan et al.’s protocol [42], based on the LM05 protocol for single photon-based QSDC, follows a similar structure with Alice acting as the certification authority. Both of these QSDC-based protocols for QIA [17, 42] are unidirectional, unlike the proposed bidirectional scheme. This highlights an advantage of the proposed protocol over in [17, 42]. Notably, this advantage extends over a range of unidirectional QIA protocols, such as Hong et al.’s protocol [50]. However, it is important to note that this advantage is not unique, as similar bidirectional capabilities are also found in Kang et al.’s protocols [19, 51] and in Zhang et al.’s 2020 work [18]. Remarkably, all these schemes [18, 19, 51] necessitate a minimum of 6 pre-shared keys for authentication, whereas our protocol requires only 4, which can be seen as advantageous due to the fact that the quantum resources are costly. Moreover, Kang et al.’s works [19, 51] utilize GHZ-like states, which are tripartite entangled states, posing challenges in preparation and maintenance compared to the Bell states utilized in our protocol. Additionally, Zhang et al.’s 2020 work assumes the third party to be semi-honest, whereas our proposed scheme considers the third party as untrusted, enhancing security. Jiang et al. proposed a semi-quantum QIA scheme resembling the ping-pong protocol [52], where Bob possesses only classical capabilities with limited quantum access, while Alice has more extensive quantum resources to prepare and measure Bell states, among other quantum operations. However, this setup introduces increased noise probability due to two-way communication and full access for Eve over the Bell state in the quantum channel. In contrast, the proposed scheme utilizes a one-way quantum channel and restricts Eve from accessing the complete Bell state. Another endeavor in QIA with QKA utilizing Bell and GHZ states with semi-honest third-party involvement was presented by Wu et al. [53]. However, their scheme demands significant quantum resources and poses maintenance challenges compared to our proposed scheme. Furthermore, their use of hash functions for authentication may not align with quantum security principles. Similarly, Li et al. proposed a simultaneous QIA scheme using GHZ states [25], but maintaining such quantum states proves difficult compared to Bell states. Thus, our proposed scheme demonstrates the desired QIA features while optimizing resource utilization, particularly when compared to the existing entangled state-based QIA schemes. The comparative analysis presented herein aims to underscore the relevance and advantages of the proposed protocol. A summarized comparison is provided in Table 2 for clarity.
Protocol Quantum resources Minimum secret key required Way of authentication Nature of the third party Secure against attacks
Zhang et al. [17] Bell state 3 Unidirectional No Impersonated fraudulent, direct measurement on channel particles, attack on two-way channel
Yuan et al. [42] Single photon 1 Unidirectional No Intercept-resend, measure-resend, entangle-measure on two-way channel
Hong et al. [50] Single photon 1 Unidirectional No Impersonation, measure resend, entangle-measure
Kang et al. [19, 51] GHZ-like 6 Bidirectional Untrusted --
Zhang et al. [18] Bell state 6 Bidirectional Semi-honest Impersonation, entangle and measure, intercept-resend, third party’s
Jiang et al. [52] Bell state $18,35$ Bidirectional No Impersonation, intercept-measure-resend, entangle-measure
Wu et al. [53] Bell state, GHZ state 6 Unidirectional Semi-honest External, dishonest participants’, third party’s, impersonation
Li et al. [25] GHZ state $10,3$ Bidirectional Trusted Impersonation, entangle-measure, intercept-measure-resend, external
Our protocol Bell state 4 Bidirectional Untrusted Impersonation, intercept-resend, impersonated fraudulent/entangle-measure
Table 2: Detailed comparison with various previous QIA protocols. The column “Minimum secret key required” contains multiple values corresponding to different protocols within the same paper.
V Conclusion
In an earlier work of the present authors [22], QIA schemes were categorized based on the intrinsic quantum cryptographic tasks involved in their design. It was observed that numerous existing schemes for secure direct quantum communication and other cryptographic tasks had been adapted to create schemes for QIA. However, the potential for modifying all types of quantum cryptographic schemes for this purpose remains largely unexplored. For instance, while some schemes for QSDC and DSQC had been adapted for QIA, the possibility of modifying their controlled versions or utilizing device-independent schemes for QSDC, and CDSQC had not been investigated. To address this gap, we present a novel QIA scheme by adapting a CDSQC scheme. In our protocol, legitimate parties use Bell states as the quantum resource and Pauli operations. With the help of an untrusted intermediary, Charlie, they authenticate each other simultaneously. We compared our protocol with previous protocols, demonstrating that our scheme achieves bidirectional authentication using fewer quantum resources than some other protocols with the help of an untrusted third party.
The present work, along with references [22, 23], suggests the potential for developing a wide range of new QIA schemes. A comprehensive exploration in this direction can be highly valuable for identifying the most efficient scheme for QIA that can be implemented using current technology. Our proposed protocol for QIA is an entangled-state-based scheme for QIA that utilizes Bell states. Inspired by the principles of secure direct quantum communication, it enables Alice to securely transmit information to Bob when controller Charlie authorizes it, without the need for pre-generated keys. Although the protocol relies solely on Bell states, its realization would necessitate quantum memory, which is not yet commercially available. While this represents a limitation, it is not unique to our protocol, but this limitation is common to many existing protocols for QIA, QSDC, DSQC and quantum dialogue. Given the growing interest in research and technology development related to quantum memory, along with recent proposals for constructing quantum memory, it is anticipated that quantum memory will become available in the near future [54]. In the interim period, a delay mechanism can serve as a substitute for quantum memory in implementing our proposed QIA protocol, thereby enhancing the resilience of protocols for secure quantum communication.
Availability of data and materials
No additional data is needed for this work.
Competing interests
The authors declare that they have no competing interests.

### v2

@@
 In the following sections, we will introduce a Bell state-based protocol for QIA, drawing from CDSQC concepts, and demonstrate its resilience against common attacks.
+Additionally, we will analyze the impact of collective noise on the proposed scheme.
 The remainder of this letter is structured as follows: Section II introduces the proposed QIA protocol that uses Bell states.
 Section III critically analyzes the protocol’s security, demonstrating its robustness against multiple types of attacks.
-Section IV compares our protocol with the existing ones from the same family.
-Finally, Section V concludes the letter.
+Section IV focuses on collective noise analysis for the new QIA scheme.
+Section V compares our protocol with the existing ones from the same family.
+Finally, Section VI concludes the letter.
 II New quantum identity authentication protocol
@@
 Moreover, an impersonated fraudulent attack entails an external party, such as Eve, aiming to mimic the legitimate user, either Alice or Bob, and effectively maneuver through the authentication procedure.
+In particular, among the attacks discussed in this section, the intercept-resend attack and impersonation-based fraudulent attack specifically target the quantum channel.
 III.1 Security assessment of the protocol against impersonation attack by Eve
@@
 In this quantum protocol, Alice and Bob generate Bell states based on their pre-shared secret key.
-Instead of transmitting transmitting their entire quantum states, Alice and Bob selectively send Particle 2 and Particle 4 through the quantum channel.
+Instead of transmitting their entire quantum states, Alice and Bob selectively send Particle 2 and Particle 4 through the quantum channel.
 This approach limits Eve’s potential attacks to these specific particles at any given moment.
@@
 3.
+The maximally mixed state is considered with equal probabilities of Bell states in Eq.
+(7) to minimize or eliminate Eve’s information gain from the channel in IR attack.
+If these probabilities are not equal, the security against the IR attack will depend on the specific values of the Bell state probabilities in the mixed state.
+As long as the mutual information between Alice and Bob remains greater than the Holevo information $\left(I(A|B)>\chi\left(\rho\right)\right)$, the scheme will remain secure against the IR attack [50].
 Figure 3: Eve’s IR attack strategy. Here, the numbers 1, 2, 3 and 4 represent the qubits in sequences $S_{A1}$, $S_{A2}$, $S_{B3}$ and $S_{B4}$, respectively.
@@
 Based on the discussion, it can be deduced that the protocol is secure against impersonated fraudulent attack by Eve, particularly under optimal conditions for Eve.
-IV Comparison with the existing protocols
+IV Collective noise analysis of proposed protocol
+The QIA protocol mentioned earlier was discussed assuming an ideal quantum channel.
+In practice, when particles are transmitted through a quantum channel, they are exposed to noise, which can alter their state(s).
+Additionally, an attacker might exploit the presence of noise to mask his attack(s), making it difficult to distinguish between errors caused by noise and those introduced by an attacker.
+Addressing the impact of collective noise on quantum communication is a well-known challenge [51, 52, 53].
+Walton et al.
+[54] introduced the concept of a decoherence-free subspace (DFS), which can mitigate the effects of collective noise since the states of the particles remain unaffected within these channels.
+Further research has also been conducted on quantum communication under collective noise [55, 56].
+Collective noise typically includes both collective-dephasing and collective-rotation noise [57].
+In what follows, we will analyze the impact of these types of noise on the scheme proposed here.
+IV.1 Collective-dephasing noise
+A collective-dephasing noise can be expressed as [57, 56]
+$\begin{array}[]{lcl}U_{dp}|0\rangle=|0\rangle&,&U_{dp}|1\rangle=e^{i\phi}|1\rangle\end{array}.$
+Collective-dephasing noise is characterized by a parameter $\phi$, which varies with time. Typically, a logical qubit encoded in the product states of two physical qubits as
+$\begin{array}[]{lclc}|0\rangle_{L}=|01\rangle&,&|1\rangle_{L}=|10\rangle&,\end{array}$
+is resistant to this type of noise because both logical qubits accumulate the same phase factor $e^{i\phi}$.
+In this context, the subscript $L$ denotes the logical qubit, while 0 and 1 correspond to horizontal and vertical polarization states, respectively; these being the eigenstates of the Pauli operator $\sigma_{z}$ ($Z$ basis).
+The states $|+_{L}\rangle$ and $|-_{L}\rangle$ are defined as follows:
+$\begin{array}[]{ccccc}|+_{L}\rangle&=&\frac{1}{\sqrt{2}}\left(|0_{L}\rangle+|1_{L}\rangle\right)&=&\frac{1}{\sqrt{2}}\left(|01\rangle+|10\rangle\right)\\ |-_{L}\rangle&=&\frac{1}{\sqrt{2}}\left(|0_{L}\rangle-|1_{L}\rangle\right)&=&\frac{1}{\sqrt{2}}\left(|01\rangle+|10\rangle\right)\end{array}.$
+From the equations presented in Section II.1, specifically from (3) and (6), it is evident that particles 2 and 4 in the composite states are responsible for transmitting quantum information through the quantum channel.
+Taking Eq.
+(3) as an example:
+$\begin{array}[]{lcl}\left|\Psi^{11}\right\rangle&=&\sigma_{z1}\otimes\sigma_{z3}\left(\left|\phi^{+}\right\rangle_{12}\left|\psi^{-}\right\rangle_{34}\right)\\ &=&\frac{1}{2}\sigma_{z1}\otimes\sigma_{z3}\left(|0001\rangle-|0010\rangle+|1101\rangle-|1110\rangle\right)_{1234}\\ &=&\frac{1}{2}\left(|0001\rangle+|0010\rangle-|1101\rangle-|1110\rangle\right)_{1234}\end{array},$
+after considering collective-dephasing noise on particles $2$ and $4$, we obtain
+$\begin{array}[]{lcl}\left|\Psi^{11}\right\rangle_{dp}&=&\frac{1}{2}\left(e^{i\phi}|0001\rangle+|0010\rangle-e^{2i\phi}|1101\rangle-e^{i\phi}|1110\rangle\right)_{1234}\\ &=&\frac{1}{2}\left(e^{i\phi}|01\rangle|00\rangle+|00\rangle|01\rangle-e^{2i\phi}|11\rangle|10\rangle-e^{i\phi}|10\rangle|11\rangle\right)_{1423}\\ &=&\frac{1}{4}\left[2e^{i\phi}\left(|\psi^{+}\rangle|\phi^{-}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle\right)+\left(1-e^{2i\phi}\right)\left(|\phi^{+}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\psi^{-}\rangle\right)_{1423}\right..\\ &+&\left.\left(1+e^{2i\phi}\right)\left(|\phi^{+}\rangle|\psi^{-}\rangle+|\phi^{-}\rangle|\psi^{+}\rangle\right)\right]_{1423}\end{array}$
+[…]
+Figure 4: The collective error probability with respect to the noise parameter can be categorized as: (a) Collective dephasing error probability with noise parameter $\phi$, and (b) Collective rotation error probability with noise parameter $\theta$.
+IV.2 Collective-rotation noise
+The collective-rotation noise can be represented as:
+$\begin{array}[]{lcl}U_{r}|0\rangle=\cos\theta|0\rangle+\sin\theta|1\rangle&,&U_{r}|1\rangle=-\sin\theta|0\rangle+\cos\theta|1\rangle\end{array},$
+The collective rotation noise is represented by the parameter $\theta$, which varies over time in the quantum channel.
+For example, applying collective rotation noise to particles 2 and 4 from Eq.
+(3) results in the following expression after a tedious calculations,
+$\begin{array}[]{lcl}U_{r2}\otimes U_{r4}\left|\Psi^{11}\right\rangle&=&\frac{1}{2}\left[-\sin\theta\cos\theta|0000\rangle-\sin^{2}\theta|0010\rangle+\cos^{2}\theta|0100\rangle+\sin\theta\cos\theta|0110\rangle\right.\\ &+&\sin\theta\cos\theta|1001\rangle-\cos^{2}\theta|1011\rangle+\sin^{2}\theta|1101\rangle-\sin\theta\cos\theta|1111\rangle\\ &-&\left(\cos^{2}\theta|0001\rangle+\sin\theta\cos\theta|0011\rangle+\sin\theta\cos\theta|0101\rangle+\sin^{2}\theta|0111\rangle\right.\\ &-&\left.\left.\sin^{2}\theta|1000\rangle+\sin\theta\cos\theta|1010\rangle+\sin\theta\cos\theta|1100\rangle-\cos^{2}\theta|1110\rangle\right)\right]_{1423}.\\ &=&\frac{1}{2}\left[2\sin\theta\cos\theta\left(-|\phi^{+}\rangle|\phi^{+}\rangle-|\psi^{-}\rangle|\psi^{-}\rangle\right)\right.\\ &+&\sin^{2}\theta\left(|\phi^{+}\rangle|\psi^{-}\rangle-|\phi^{-}\rangle|\psi^{+}\rangle+|\psi^{+}\rangle|\phi^{-}\rangle-|\psi^{-}\rangle|\phi^{+}\rangle\right)\\ &+&\left.\cos^{2}\theta\left(|\psi^{+}\rangle|\phi^{-}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle-|\phi^{+}\rangle|\psi^{-}\rangle-|\phi^{-}\rangle|\psi^{+}\rangle\right)\right]_{1423}\end{array}$
+In the absence of noise or errors, the Bell pairs shared by Alice and Bob are $\left\{|\psi^{-}\rangle|\phi^{+}\rangle,|\psi^{+}\rangle|\phi^{-}\rangle,|\phi^{-}\rangle|\psi^{+}\rangle,|\phi^{+}\rangle|\psi^{-}\rangle\right\}$.
+From this above result, the probability of error due to collective rotation noise can be derived as $2\sin^{2}\theta\cos^{2}\theta$, which depends on the noise parameter $\theta$.
+Similarly, applying collective rotation noise to particles 2 and 4 from Eq.
+(4) yields:
+$\begin{array}[]{lcl}U_{r2}\otimes U_{r4}\left|\Psi^{00}\right\rangle&=&\frac{1}{2}\left[\cos^{2}\theta|0000\rangle+\sin\theta\cos\theta|0010\rangle+\sin\theta\cos\theta|0100\rangle+\sin^{2}\theta|0110\rangle\right.\\ &+&\sin^{2}\theta|1001\rangle-\sin\theta\cos\theta|1011\rangle-\sin\theta\cos\theta|1101\rangle+\cos^{2}\theta|1111\rangle\\ &-&\left(-\sin\theta\cos\theta|0001\rangle-\sin^{2}\theta|0011\rangle+\cos^{2}\theta|0101\rangle+\sin\theta\cos\theta|0111\rangle\right.\\ &-&\left.\left.-\sin\theta\cos\theta|1000\rangle+\cos^{2}\theta|1010\rangle-\sin^{2}\theta|1100\rangle+-\sin\theta\cos\theta|1110\rangle\right)\right]_{1423}.\\ &=&\frac{1}{2}\left[2\sin\theta\cos\theta\left(|\phi^{-}\rangle|\psi^{+}\rangle+|\psi^{+}\rangle|\phi^{-}\rangle\right)\right.\\ &+&\sin^{2}\theta\left(|\psi^{+}\rangle|\psi^{+}\rangle-|\psi^{-}\rangle|\psi^{-}\rangle+|\phi^{+}\rangle|\phi^{+}\rangle-|\phi^{-}\rangle|\phi^{-}\rangle\right)\\ &+&\left.\cos^{2}\theta\left(|\phi^{+}\rangle|\phi^{+}+|\phi^{-}\rangle|\phi^{-}\rangle+|\psi^{+}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\psi^{-}\rangle\right)\right]_{1423}\end{array}$
+[…]
+V Comparison with the existing protocols
 In this section, we aim to conduct a concise comparative analysis involving the proposed protocol and a selection of previously proposed QIA schemes.
@@
 Table 2: Detailed comparison with various previous QIA protocols. The column “Minimum secret key required” contains multiple values corresponding to different protocols within the same paper.
-V Conclusion
+VI Conclusion
 In an earlier work of the present authors [22], QIA schemes were categorized based on the intrinsic quantum cryptographic tasks involved in their design.
@@
 In the interim period, a delay mechanism can serve as a substitute for quantum memory in implementing our proposed QIA protocol, thereby enhancing the resilience of protocols for secure quantum communication.
+Additionally, our work could be extended in the future with modifications that provide resistance to collective noise [58, 59, 60].
 Availability of data and materials